# The Fraction of Clumpy Galaxies in JWST Surveys over $2<z<12$

### v1

@@ 3.2 The Final Sample
Note. — Each fraction is in terms of the rest-frame NUV luminosity of the host galaxy. Empty values indicate that either too few (<10<10) galaxies lie in a given mass and redshift bin or clumps could not be detected with at least 50% completeness.
4 Selection of Clumpy Galaxies
4.1 Identification of Clumps
Clumps are identified for all galaxies in the mass-selected, star-forming sample (Section 3.2) using the technique developed by Sattari et al. (2023). Briefly, it involves taking the Fourier Transform of the rest-frame NUV image of a galaxy, filtering out the low-frequency signal that corresponds to the host galaxy light, and performing source detection on the leftover high-frequency signal, which consists of clumps and noise (see Sec. 3 in Sattari et al. 2023 for details). To distinguish these, we apply stringent source detection criteria.
Objects are identified as clumps in JWST/NIRCam images if they span at least five contiguous pixels, each of which must be at least two standard deviations brighter than the background-subtracted image, and if the total S/N of the object is at least 3. These criteria are the same as those used by Shibuya et al. (2016), who used HST data, and similar to those used by Claeyssens et al. (2025), who used JWST data. We use photutils to detect clumps. For a given galaxy, clumps must lie at least 0.″1 from the center and within the segmentation map created by the JADES team. This corresponds to a distance of 0.86 kpc at $z=2$ and 0.37 kpc at $z=12$. The former criterion is used to avoid confusion with a bulge or active galactic nucleus and the latter ensures that all detected clumps lie within a single galaxy and not within its neighbors projected on the sky.
Figure 6: Example clumpy galaxies are shown in this figure for each of the stellar mass and redshift bins considered in this work (Figure 5 and Section 4.2). Stellar mass bins are indicated with the columns and increase from left to right. Redshift bins are denoted with the rows and increase from top to bottom. For each galaxy in a given mass and redshift bin, images in the rest-frame NUV are displayed in the left two columns and a RGB image is shown in the rightmost column. Images are shown with an arcsinh stretch. In the middle columns, detected clumps are circled in pink. Clumps that meet the completeness limits listed in Table 1 are denoted with solid lines and those that do not with dotted lines. The segmentation map of each galaxy is indicated with the red outline. The bandpasses used to create the RGB images are listed on the far left side of the figure.
4.2 Completeness of Clump Identification
To ensure that a complete sample of clumps is identified, we measure how faint clumps can be in order to detect at least half of them, following pre-JWST studies (Guo et al., 2015; Shibuya et al., 2016; Sattari et al., 2023). This is done by adding fake clumps to images of randomly selected real galaxies, regardless whether they have clumps, and running the clump identification method described above.
Fake clumps are modeled as point sources following previous works (e.g., Guo et al., 2012, 2015, 2018; Shibuya et al., 2016; Soto et al., 2017; Sattari et al., 2023). We measure empirical point spread functions (PSFs) for all NIRCam bandpasses in GOODS-N and GOODS-S. This is described in Appendix A. Positions of the fake clumps are randomly selected to lie within the segmentation map of a given galaxy. Their brightnesses vary randomly between 0.1% and 75% of that of the host galaxy. In order to have sufficient statistics to determine the completeness, 100 galaxies are randomly selected in each of three stellar mass bins ($8<\log\left(M_{\star}/M_{\odot}\right)<9,9<\log\left(M_{\star}/M_{\odot}\right)<10$, and $10<\log\left(M_{\star}/M_{\odot}\right)<11$) for each redshift bin. If fewer than 100 galaxies exist within a given bin in mass and redshift, all galaxies within that bin are selected. In total, 1,541 galaxies are randomly selected. A fake clump is added to the rest-frame NUV image of each galaxy using the ImagePSF routine in photutils. The clump identification algorithm is then run on the modified image to determine whether the fake clump is recovered. For each randomly selected galaxy, this procedure is repeated 50 times. The position and brightness of the fake clumps are randomly varied in each iteration. Uncertainties on the completeness are computed in a Bayesian fashion and assume a noninformative Jeffreys prior on the binomial distribution (Cameron, 2011).
The completeness of the fake clump simulations is shown in Figure 5. In each panel, the fraction of clumps detected by the clump identification algorithm is shown as a function of the fake clump luminosity. This is further shown as a function of stellar mass (lines of various thicknesses) and redshift (columns). Clumps are more likely to be detected in higher-mass systems and lower redshifts. The minimum fractional luminosities at which clumps are detected at 50% completeness are indicated with the vertical dashed-dotted lines in Figure 5. For galaxies at $z\sim 2.75$ and over $9<\log\left(M_{\star}/M_{\odot}\right)<10$, 50% of clumps are detected when they are at least 2% of the NUV luminosities of their hosts. At $z\sim 7.75$ and over the same range in mass, the same completeness is obtained for a minimum fractional luminosity of 7%. At lower masses, minimum fractional luminosities are higher, and at higher masses, they are similar. At $z\sim 10.5$, no clumps are detected with at least 50% completeness. Because of this, galaxies in the two highest redshift bins at $z\sim 7.75$ and $z\sim 10.5$ will be combined into one redshift bin in the following Sections. The minimum fractional luminosities at $z\sim 7.75$ will be used to find clumps over $6.5<z<12$.
Table 1 lists the minimum fractional luminosities as a function of stellar mass and redshift. At $z\sim 2.75$ and $\log\left(M_{\star}/M_{\odot}\right)>9$, clump identification is 50% complete at four times fainter luminosities with JWST (2%) than with HST (8%; Shibuya et al. 2016).
A galaxy is defined to be clumpy if it has at least one off-center clump with a rest-frame NUV luminosity that is greater than or equal to the completeness limits listed in Table 1. These depend on stellar mass and redshift. Example clumpy galaxies are shown in Figure 6. Clumps that are brighter than the completeness limits are indicated with solid circles and those that are fainter with dotted circles.
Figure 7: The fraction of clumpy galaxies, $f_{\rm{clumpy}}$, over $2<z<12$. It peaks at $z\sim 2.75$ and declines by a factor of $\sim 4$ from $z\sim 2.75$ to $z>6.5$ at a nearly constant rate. Galaxies are considered clumpy if they have at least one off-center clump. Measurements from this work for galaxies with stellar masses of at least $10^{9}~M_{\odot}$ are shown as black hexagons connected by a solid black line. $f_{\rm{clumpy}}$ measurements for galaxies in GOODS-N are shown to the left of the solid black line. Those for galaxies in GOODS-S are to the right of it. The minimum fractional luminosity changes with redshift according to Table 1. We compile measurements from the literature and plot them alongside ours. These are shown as various symbols and lines and are listed in Table 2 (see Section 5 for details). The results are consistent: $f_{\rm{clumpy}}$ rises gradually from $z\sim 8$ to $z\sim 1.5-2$, when it peaks, and declines at lower redshifts. This is qualitatively consistent with the evolution of the cosmic SFR density, which is shown as a thin blue solid line and has been arbitrarily scaled vertically. The results of this study are consistent with the view that clumpy galaxies became more common at later times. However, the $f_{\rm{clumpy}}$ calculated in this work is higher at fixed redshift and increases more steeply with decreasing redshift than the bulk of previous studies. This suggests that clumpy galaxies were more common at high redshift than previously measured, but with consistent trends towards lower redshifts.
4.3 Aperture Photometry of Clumps
We measure the fluxes of detected clumps by performing fixed aperture photometry with photutils. Each aperture is circular with a radius of 3 pixels. The background is measured in annuli with inner radii of 4 pixels and outer radii of 10 pixels. To perform background subtraction, the mean surface brightness of the light in the annuli is measured, multiplied by the area of the aperture, and subtracted. In each of these steps, for a given galaxy, any nearby clumps are masked using a circular mask of radius equal to 4 pixels. The sizes of the apertures, background annuli, and masks are adopted from Guo et al. (2018), who carried out extensive tests on photometry of clumps in HST images.
Because clumps are modeled as point sources, the circular apertures used here enclose the majority of the total light, about 50-70%, depending on the bandpass. We account for this by applying aperture corrections that vary with wavelength. Clump aperture fluxes are multiplied by factors of 1.434, 1.400, 1.422, 1.477, and 1.886 in GOODS-N and 1.471, 1.429, 1.469, 1.502, and 1.806 in GOODS-S for the F090W, F115W, F150W, F200W, and F277W bandpasses, respectively. These are determined using the empirical PSFs that are measured in this work (see Appendix A).
5 The Fraction of Clumpy Galaxies over the Redshift Range $2<z<12$
We now study the fraction of clumpy galaxies ($f_{\rm{clumpy}}$) as a function of redshift. For a galaxy to be considered clumpy, it must have at least one off-center clump identified in its rest-frame NUV image. This is done using the technique developed by Sattari et al. (2023). Clumps are identified using the minimum fractional luminosities listed in Table 1). The fraction of clumpy galaxies is defined to be the number of galaxies with clumps divided by the total number of galaxies in a given bin in stellar mass, redshift, and/or sSFR. All uncertainties are calculated assuming Bayesian confidence intervals on the binomial distribution, unless otherwise noted.
The evolution of $f_{\rm{clumpy}}$ with redshift is examined in Section 5.1. In Section 5.2, measurements from this work are compared with those from other studies that overlap with ours in redshift. We find that our results disagree with those from most studies over $2<z<6.5$. To understand why this is, the redshift evolution of $f_{\rm{clumpy}}$ is measured from HST and JWST images and examined in Section 5.3.
Table 2: Summary of Previous Studies on Clumpy Galaxies
@@ 5 The Fraction of Clumpy Galaxies over the Redshift Range $2<z<12$
Conselice & Arnold (2009) LBGs (133) $\gtrsim 9$ 4–6 $CASGM_{20}$eefootnotemark: rest-frame UV
@@
Tan et al. (2024) MWAfffootnotemark: progenitors (909) 8.5–10.7 0.25–5 Algorithm rest-frame $U$ band
@@
This work SFGs (9121) $\geq 8$ 2–12 Algorithm rest-frame NUV
@@
Note. — aLyman-break Galaxies. bStellar mass ranges for samples comprising Lyman-break galaxies are estimated using the relations between absolute UV magnitude and stellar mass shown in Fig. 1 in Shibuya et al. (2015). cC:C:”concentration; “A:A:”asymmetry; “G:G:” Gini coefficient; $M_{20}:$ second-order moment parameter. dStar-forming galaxies. eS:S:” clumpiness. fMilky Way Analog. See Conselice (2014) for definitions of the $CASGM_{20}$ morphological parameters.
5.1 Evolution with Redshift
@@
Figure 7 shows $f_{\rm{clumpy}}$ as a function of redshift. Measurements from this work for galaxies with stellar masses greater than $10^{9}~M_{\odot}$ are shown as black hexagons connected by a thick solid line. This limit in mass is constant with redshift and chosen for two reasons. First, at a fixed redshift, clumps are identified using a uniform cut in luminosity at masses above this limit (see Table 1). Second, it is comparable to the mass limits used by other studies (see Table 2).
@@
The fraction of clumpy galaxies is highest ($\sim 80\%$) at $z\sim 2.75$ and decreases with increasing redshift to $\sim 20$% at $z>6.5$. The fraction decreases at a roughly constant rate with increasing redshift: it is $\sim 60\%$ at $z\sim 4.25$ and $\sim 40\%$ at $z\sim 5.75$. $f_{\rm{clumpy}}$ measurements in either GOODS-N or GOODS-S agree with this trend and differ from it by $\leq 4\%$. These are indicated with empty hexagons to the left and right of the solid black line in Figure 7.
@@
Our $f_{\rm{clumpy}}$ measurements are also compared with the cosmic SFR density (SFRD), as measured by Madau & Dickinson (2014). This is shown as a thin blue solid line in Figure 7. The fraction of clumpy galaxies is observed to rise with decreasing redshift more steeply than the SFRD. However, the net change in $f_{\rm{clumpy}}$ from $z>6.5$ to $z\sim 2.75$, an increase of about a factor of four, is similar to that in the SFRD over the same range in redshift. This seemingly suggests a connection between the frequency of clumpy galaxies and the volume-averaged SFR at high redshift.
Figure 8: Clumpy galaxies are found to be more common over $2<z<6$ in this work compared with earlier studies based on HST due to the higher sensitivity of JWST. This figure shows $f_{\rm{clumpy}}$ measured with HST (dashed lines) and JWST (solid lines) as a function of redshift for three choices of the minimum fractional luminosity $(L_{\rm{clump}}/L_{\rm{host~galaxy}})$. This is the minimum fraction of the host galaxy NUV luminosity a clump must have. A slight offset in redshift is applied to each set of $f_{\rm{clumpy}}$ measurements for clarity. In the left panel, the limit is set to 8%, which is adopted from pre-JWST studies. Measurements from both telescopes agree to within 9% and are statistically consistent. At the intermediate limit of 5%, shown in the middle panel, fractions differ by 6–18%, with those from JWST being higher at all redshifts. Estimates are statistically consistent for only one redshift bin. Finally, at the lowest limit of 2%, shown in the right panel, JWST-based measurements are higher than those from HST by 22–29% and statistically inconsistent at all redshifts.
5.2 Comparison with Other Studies
@@
Our $f_{\rm{clumpy}}$ measurements are also compared with those from other studies in Figure 7. The samples and methods of these are summarized in Table 2. The uncertainties on $f_{\rm{clumpy}}$ were not published in some studies. They were calculated in this work assuming Bayesian binomial confidence intervals, wherever feasible. Details on the extraction of $f_{\rm{clumpy}}$ measurements from earlier studies are provided in Appendix B.
@@
We now examine the evolution of $f_{\rm{clumpy}}$ over $0<z<8$ in the rest-frame UV from previous studies, shown in Figure 7. This was found to be $\sim 30-40\%$ at $z\sim 0.5$, peak at $\sim 60-80\%$ at $z\sim 2$, and decline to $\sim 40-60\%$ at $z\sim 3$ (Elmegreen et al. 2007; Wuyts et al. 2012; Tadaki et al. 2014; Guo et al. 2015; Shibuya et al. 2016; Sattari et al. 2023; although lower fractions of $\sim 20-40\%$ were obtained by Law et al. 2012). The decline continues to $\sim 10-30\%$ at $z\sim 4$ (Lotz et al. 2006; Ravindranath et al. 2006; Conselice & Arnold 2009; Shibuya et al. 2016; Ribeiro et al. 2017; although Elmegreen et al. 2007 and Curtis-Lake et al. 2016 measured higher fractions of $\sim 40-50\%$). At $z\gtrsim 4$, previous measurements of $f_{\rm{clumpy}}$ largely divide into two groups: lower fractions of $\sim 10-20\%$ that are approximately constant with redshift (Conselice & Arnold 2009; Oesch et al. 2010; Shibuya et al. 2016; and Kawamata et al. 2015, although their $f_{\rm{clumpy}}$ was measured to be $\sim 50$% at $z\sim 8$) and higher fractions of $\sim 30-70\%$ over $4<z<8$ (Elmegreen et al., 2007; Jiang et al., 2013; Curtis-Lake et al., 2016; Bowler et al., 2017; Ribeiro et al., 2017; Harikane et al., 2025). Another measurement of $f_{\rm{clumpy}}$ as a function of redshift was recently made by Tan et al. (2024). Galaxies were selected in bins of stellar mass that evolve with redshift to track the redshift evolution of progenitors of Milky Way-like systems. Their evolution of $f_{\rm{clumpy}}$ with redshift is consistent with the bulk of studies discussed here: $f_{\rm{clumpy}}$ was measured to be $\sim 20\%$ over $2<z<5$ and increased with redshift to $\sim 60\%$ at $z\sim 0.5$.
The results of this work over $2<z<6.5$ are inconsistent with those from most of the previous studies at similar redshifts. We find that $f_{\rm{clumpy}}$ rose quickly with redshift at early times, from $\sim 20\%$ at $z>6.5$ to $\sim 80\%$ at $z\sim 2.75$ at a roughly constant rate with redshift. In contrast, most studies found that $f_{\rm{clumpy}}$ changed by $\lesssim 10\%$ over $4<z<8$ and began to increase at $z\lesssim 4$.
Our measurements are consistent with those from Wuyts et al. (2012), Jiang et al. (2013), and Bowler et al. (2017) at the redshifts where they overlap. Our results are also consistent with those from Elmegreen et al. (2007) and Ribeiro et al. (2017) at $z\sim 4.25$ and those from Shibuya et al. (2016) at $z>6.5$.
5.3 Comparing Clumpy Galaxy Fractions between the HST and JWST Data
@@
Our results on the redshift evolution of $f_{\rm{clumpy}}$ disagree with those of most studies over $2<z<6.5$. All but two of these are based on HST data, whereas this work is based on JWST data. To determine the cause of the disagreement, we compute $f_{\rm{clumpy}}$ using both HST and JWST images for a subset of our sample and vary the minimum fractional luminosity of the clumps. As we demonstrate in this Section, it is largely the difference in sensitivity between the two datasets that is responsible for the disagreement in the $f_{\rm{clumpy}}$ measurements shown here. Other factors may have secondary effects on the measurement of $f_{\rm{clumpy}}$, such as the differences between photometric redshifts, stellar masses, and SFRs from earlier studies and those used in this work.
@@
Galaxies were selected for this comparison as follows. We match the mass-selected, star-forming sample (Section 3.2) with photometric catalogs from the CANDELS survey, which were created by Barro et al. (2019) for GOODS-N and Guo et al. (2013) for GOODS-S. We use a match radius of 0.″5 and match HST/Wide Field Camera 3 (WFC3) F160W magnitudes to within 0.5 mag. There are 2,007 matches in GOODS-N and 2,189 in GOODS-S. As studies based on HST data limited their analyses to galaxies with stellar masses $\gtrsim 10^{9}~M_{\odot}$ (see Table 2), we only consider those with masses $\geq 10^{9}~M_{\odot}$. This mass cut leaves a total of 1,245 galaxies (592 in GOODS-N and 653 in GOODS-S) that will be used for our comparison.

### v2

@@ V Fraction of Clumpy Galaxies and Their Evolution with Redshift
Note. — Each fraction is in terms of the rest-frame NUV luminosity of the host galaxy. Empty values indicate that either too few (<10<10) galaxies lie in a given mass and redshift bin or clumps could not be detected with at least 50% completeness. All minimum fractional luminosities here are reported for the optimal smoothing length of 10 pixels.
Table 2: Summary of Literature on Clumpy Galaxies
@@
Conselice and Arnold (2009) LBGs (133) $\gtrsim 9$ 4–6 $CASGM_{20}$eefootnotemark: rest-frame UV
@@
This work SFGs (9121) $\geq 8$ 2– 9 Algorithm rest-frame NUV
@@
Note. — aLyman-break Galaxies. bStellar mass ranges for samples comprising Lyman-break galaxies are estimated using the relations between absolute UV magnitude and stellar mass shown in Fig. 1 in Shibuya et al. (2015). cC:C:”concentration; “A:A:”asymmetry; “G:G:” Gini coefficient; $M_{20}:$ second-order moment parameter. dStar-forming galaxies. eS:S:” clumpiness. See Conselice (2014) for definitions of the $CASGM_{20}$ morphological parameters.
V.1 Evolution with Redshift
@@
Figure 8 shows $f_{\rm{clumpy}}$ as a function of redshift. Measurements from this work for galaxies with stellar masses greater than $10^{9}~M_{\odot}$ are shown as black hexagons connected by a thick solid line. This limit in mass is constant with redshift and chosen because it is comparable to the mass limits used by other studies (see Table 2).
@@
The fraction of clumpy galaxies is highest ($\sim 70\%$) at $z\sim 2.75$ and decreases with increasing redshift to $\sim 20$% at $z\sim 7.75$. It changes the most from $\sim 30\%$ at $z\sim 5.75$ to $\sim 60\%$ at $z\sim 4.25$. $f_{\rm{clumpy}}$ measurements in either GOODS-N or GOODS-S agree with this trend within $\leq 3\%$.
@@
Our $f_{\rm{clumpy}}$ measurements are also compared with the cosmic SFR density (SFRD), as measured by Madau and Dickinson (2014), in Figure 8. It has been rescaled along the y-axis to line up with the majority of the points in the figure. The fraction of clumpy galaxies rise s with decreasing redshift more steeply than the SFRD. However, the net change in $f_{\rm{clumpy}}$ from $z\sim 7.75$} to $z\sim 2.75$, an increase of almost a factor of four, is similar to that in the SFRD over the same range in redshift. This seemingly suggests a connection between the frequency of clumpy galaxies and the volume-averaged SFR as a function of redshift.
V.2 Comparison with Other Studies
@@
Our $f_{\rm{clumpy}}$ measurements are also compared with those from other studies in Figure 8. The samples and methods of these are summarized in Table 2. The uncertainties on $f_{\rm{clumpy}}$ were not published in some studies. They were calculated in this work assuming Bayesian binomial confidence intervals, wherever feasible. Details on the extraction of $f_{\rm{clumpy}}$ measurements from earlier studies are provided below.
Following Shibuya et al. (2016), we include $f_{\rm{clumpy}}$ measurements from Lotz et al. (2006); Ravindranath et al. (2006); Conselice and Arnold (2009); Oesch et al. (2010); Law et al. (2012); Jiang et al. (2013); Kawamata et al. (2015) and Curtis-Lake et al. (2016). These were calculated using the fractions of galaxies undergoing mergers and/or exhibiting disturbed morphologies or multiple cores. In Ravindranath et al. (2006), Oesch et al. (2010), and Kawamata et al. (2015), morphologies were determined via visual inspection. In the other works, morphological parameters were calculated.
Following Guo et al. (2015), we also include $f_{\rm{clumpy}}$ measurements from Elmegreen et al. (2007). In the latter work, galaxies were visually inspected and categorized. $f_{\rm{clumpy}}$ was inferred using the abundances of clump cluster, spiral, and elliptical galaxies. Other categories (chain galaxies, double nuclei, and tadpoles) were excluded as they generally have small axis ratios and would not have been included according to our sample selection (Section III).
$f_{\rm{clumpy}}$ from Bowler et al. (2017) was obtained from their merger fraction, which was determined via visual inspection.
For all other studies listed in Table 2, $f_{\rm{clumpy}}$ measurements were given explicitly. In Wuyts et al. (2012), $f_{\rm{clumpy}}$ was measured by searching for galactic regions with blue colors and high surface brightnesses, based on visual inspection of co-added surface brightness profiles. Clumps in Tadaki et al. (2014) were identified with the clumpfind code (Williams et al., 1994), which uses multiple isophotes to find clumps. In Guo et al. (2015) and Shibuya et al. (2016), clumps are found using algorithms that are nearly identical to the one described in Section IV.1. Ribeiro et al. (2017) detected clumps by identifying disconnected regions at various isophotal thresholds. Huertas-Company et al. (2020) used neural networks to detect clumps with a training set of simulated images of galaxies. Sattari et al. (2023) identified clumps by modeling the smooth, background light of host galaxies with Fourier transforms and subtracting this model to reveal clumps. Detection techniques similar to those used in Guo et al. (2015) and Shibuya et al. (2016) were then applied to the background-subtracted images. Last, clumps were detected in Harikane et al. (2025) by performing source detection on images in the F115W bandpass. Subcomponents detected with $5\sigma$ significance were then visually inspected.
@@
For the rest of the studies shown in Figure 8, we find $f_{\rm{clumpy}}$ to be $\sim 30-40\%$ at $z\sim 0.5$, peak at $\sim 60-80\%$ at $z\sim 2$, and decline to $\sim 40-60\%$ at $z\sim 3$ (although lower fractions of $\sim 20-40\%$ were obtained by Law et al. 2012). The decline continues to $\sim 10-30\%$ at $z\sim 4$ (although Elmegreen et al. 2007 and Curtis-Lake et al. 2016 measured higher fractions of $\sim 40-50\%$). At $z\gtrsim 4$, previous measurements of $f_{\rm{clumpy}}$ largely divide into two groups: lower fractions of $\sim 10-20\%$ that are approximately constant with redshift and higher fractions of $\sim 30-70\%$ over $4<z<8$.
The results of this work are consistent with those of most previous studies at $z\gtrsim 5.75$ and largely inconsistent at lower redshifts. We find that, as redshift decreases, $f_{\rm{clumpy}}$ rises more steeply in this work than in the bulk of earlier studies. At $z\gtrsim 5.75$, $f_{\rm{clumpy}}$ as measured in this work is $\sim 20\%$, similar to what was measured earlier at these redshifts. At $z\sim 4.25$, we find $f_{\rm{clumpy}}$ to be $\sim 60\%$, whereas most earlier works found values between 20 and 40%. At $z\sim 2.75$, our measure of $f_{\rm{clumpy}}$ is $\sim 70\%$, while most other works found it to be between 40 and 60%.
V.3 Comparing Clumpy Galaxy Fractions between the HST and JWST Data
Figure 9: To understand why $f_{\rm{clumpy}}$ is higher with JWST data compared with previous data from HST, it is calculated with images from both telescopes as a function of redshift in each panel, minimum fractional luminosity (a proxy for image depth, columns) and whether JWST images are at their original pixel scale and resolutions (top row) or those of HST images (bottom). $f_{\rm{clumpy}}$ from JWST data, as shown with the large hexagons and solid lines, lies systematically above that from HST data, plotted as small circles and dashed lines. As the minimum fractional luminosity gradually decreases, from left to right, $f_{\rm{clumpy}}$ from JWST data progressively increases. Importantly, this is seen in both sets of JWST images, which are consistent with each other to within 9%, at worst. However, $f_{\rm{clumpy}}$ from HST images barely changes. Taken together, this suggests that the increased depth of JWST is the primary reason for the higher $f_{\rm{clumpy}}$ measured in this work. Difference in resolution between HST and JWST plays a secondary role. Clumps are treated as point sources for only this comparison, as they were found to have light profiles consistent with point sources in work based on HST data.
@@ V.3 Comparing Clumpy Galaxy Fractions between the HST and JWST Data
Our results on the redshift evolution of $f_{\rm{clumpy}}$ disagree with those of most studies listed in Table 2 at $z\sim 2.75$ and $z\sim 4.25$. All but one of these are based on the HST data, whereas this work is entirely based on JWST data. To determine the cause of the disagreement, we compute $f_{\rm{clumpy}}$ using both HST and JWST images for a subset of our sample. We measure $f_{\rm{clumpy}}$ as a function of redshift at various minimum fractional luminosities. The aim is to assess to what extent the difference in resolution and sensitivity between the telescopes affect the measurement of $f_{\rm{clumpy}}$. In the following, we demonstrate that it is largely the difference in sensitivity that is responsible for the disagreement in the $f_{\rm{clumpy}}$ measurements, with difference in resolution having a secondary effect.
@@
To select galaxies for this comparison, we match the mass-selected, star-forming sample (end of Section III) with photometric catalogs from GOODS-N (Barro et al., 2019) and GOODS-S (Guo et al., 2013). We use a match radius of 0.″5 and match HST/Wide Field Camera 3 (WFC3) F160W magnitudes from these catalogs and those from the JADES to within 0.5 mag. There are 2,007 matches in GOODS-N and 2,189 in GOODS-S. As studies based on HST data limited their analyses to galaxies with stellar masses $\gtrsim 10^{9}~M_{\odot}$ (see Table 2), we only consider this mass range. This leaves a total of 1,245 galaxies (592 in GOODS-N and 653 in GOODS-S) for our comparison.